Suppose the demand equation for computers by Teetan Ltd for the year 2017 is given by Qd= 1200-P and the supply equation is given by Qs= 120+3P. Find equilibrium price and analyse what would be the excess demand or supply if price changes to Rs 400 and Rs 120.


1
Expert's answer
2020-12-10T09:43:31-0500

Qd=Qs

1200-P=120+3P

P=270

P=400

"Qd=1200-400=800"

"Qs=120+3*400=1320"

P=120

"Qd=1200-120=1080"

"Qs=120+3*120=480"